While spraying your lawn with 2-4D (an herbicide) on a windy day, about 25 g blows into your swimming pool (volume 500,000 L). The maximum acceptable concentration for 2-4D in drinking water is 0.10 ppm. Would you swim in the pool? Justify your answer
1
Expert's answer
2020-11-17T10:27:21-0500

Parts Per Million (ppm) is a measurement of the concentration of dilute solution.

In weight/weight (w/w) terms 1 ppm = 1 mg solute / 1 kg solution.

Convert the mass of the solute in grams to a mass in milligrams:

25 g of 2-4D × 1000 mg = 25000 mg of 2-4D

the mass of the water solution in liters is the mass in kilograms:

500000 L of solution = 500000 kg of solution

ppm of 2-4D = 25000mg / 500000kg = 0.05 ppm < 0.1 ppm

so I can swim in the pool.

Answer: Concentration of 2-4D in the pool is 0.05 ppm. I can swim in the pool.